On 1 January 2016, an optimistic energy trader purchased a 90 day Asian (arithmetic) call option on crude oil futures. The option has an exercise price of $35. The of crude oil futures at the 30th, 60th and 90th day of the option was $30.65, $36.90, and $38.49. What is the payoff to the option?
Answer #1

Arithmetic mean of stock prices = (30.65 + 36.90 + 38.49)/3 = $35.34

The profit is the average(mean) minus the strike price, which is = 35.34 - 35 = 0.34 dollars